A laser is used in eye surgery to weld a detached retina back into place. The wavelength of the laser beam is 468 nm, while the power is 1.9 W. During surgery, the laser beam is turned on for 0.070 s. During this time, how many photons are emitted by the laser?
Answer #1

Solution)

We know,

Energy of one photon = hc/λ = 6.626 x 10^-34 * 3.00 x 10^8 / 5468 x10^-9 = 3.63x 10^-17 Joules

Total energy = power * time = 1.9 *0.070 =  0.133 Joules

Number of photons = total energy / energy perphoton = 0.133 / 3.63 x10^-17= 3.66 x10^15 photons

============